# Beyond Recommendations: From Backward to Forward AI Support of Pilots’ Decision-Making Process

### v1

@@ Abstract.
journal: PACMHCIjournalvolume: 8journalnumber: CSCW2journalyear: 2024publicationmonth: 10copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Decision support systemsccs: Human-centered computing Interaction paradigms
@@
Schematic diagram of both support concepts.
@@ 1. Introduction
Yet, in spite of good machine performance, real-world adoption of AI is often difficult (Schlögl et al., 2019; Varghese, 2020). In controlled studies, failure to achieve complementary performance is often observed (Bansal et al., 2021; Liu et al., 2021; Fok and Weld, 2023), i.e., the combination of human and AI performs worse than one of them alone. This is the result of inappropriate reliance, including both overreliance (human relies on AI even when it is disadvantageous to do so) (Bansal et al., 2021; Liu et al., 2021; Jacobs et al., 2021b; Schmidt and Biessmann, 2020; Bussone et al., 2015; Buçinca et al., 2021; Lai and Tan, 2019; Poursabzi-Sangdeh et al., 2021; Wang and Yin, 2021; Green and Chen, 2019) and underreliance (human rejects AI even when it would be beneficial to rely on it) (Dietvorst et al., 2015; Cheng and Chouldechova, 2023; Castelo et al., 2019; Prahl and Van Swol, 2017).
@@
In practice, AI support often turns out less useful to decision makers than imagined (Yang et al., 2016; Blomberg et al., 2018; Kawakami et al., 2022a). A growing number of formative studies on real-world tasks with domain experts (Kawakami et al., 2022a, b; Jacobs et al., 2021a; Sivaraman et al., 2023; Bach et al., 2023; Burgess et al., 2023; Kaltenhauser et al., 2020; Yang et al., 2016; Zhang et al., 2023b) try to understand what hinders effective use of AI decision support. A frequent problem is that AI decision support is usually designed to be recommendation-centric, where the primary functionality of the system is to give end-to-end decision recommendations, i.e., the system suggests a possible end result straight from its input data. By directly jumping to the end result, these systems only support the very end of the decision-making (Buçinca et al., 2022), ignoring the entire process leading up to the decision (Yang et al., 2016; Zhang et al., 2023b, a).
@@
While the limitations of recommendation-centric support are becoming increasingly more apparent, there is a notable lack of effective solutions to the identified problems. Current research is mostly limited to proposing alternatives for recommendation-centric support on a conceptual level (Zhang et al., 2021; Buçinca et al., 2022; Cabitza et al., 2021; Miller, 2023; Koon, 2022), but concrete examples are few and far between, with most examples stemming from healthcare (Cai et al., 2019; van Berkel et al., 2021; Yang et al., 2019; Lindvall et al., 2021; Gu et al., 2023; Zhang et al., 2023b). Even fewer works evaluate these alternative support paradigms in comparison to recommendation-centric decision support.
@@
One of the few studies exploring alternative decision support roles for AI is Zhang et al.’s work on diversion assistance (Zhang et al., 2023a). Their key insight was that diversion decisions are not a point, but a process in which pilots take proactive actions. Even during normal flight, pilots constantly establish situation awareness (SA) and prepare a valid plan B, should an emergency occur. Continuously supporting these proactive actions via unobtrusive hints was found to be a promising support role for AI, but it remains unclear how effective it would be in practice, especially compared to the established recommendation-centric paradigm.
@@ 2.1. Recommendation-Centric Decision Support
The most common strategy to help decision makers work better with recommendation-centric AI is to add explanations of how the model works (Wang and Yin, 2021; Bussone et al., 2015; Bansal et al., 2021; Zhang et al., 2020; Lai and Tan, 2019) and other model information, such as model confidence (Zhang et al., 2020; McGuirl and Sarter, 2006; Rechkemmer and Yin, 2022) or the stated model accuracy (He et al., 2023; Rechkemmer and Yin, 2022). The goal is to help people to rely appropriately on AI recommendations (Schemmer et al., 2023) by giving cues about when they may be beneficial or detrimental to rely on. Results have been mixed so far, as especially explanations are prone to induce blind trust (Bansal et al., 2021; Eiband et al., 2019) and hence overreliance, even among domain experts (Bussone et al., 2015; Jacobs et al., 2021b). Recently, Vasconcelos et al. (Vasconcelos et al., 2023) have shown that explanations can reduce overreliance under certain conditions, but it is questionable how often these conditions are valid in real applications (Fok and Weld, 2023). Communicating model confidence appears more promising, as it has repeatedly been shown to improve appropriate reliance (Zhang et al., 2020; McGuirl and Sarter, 2006; Bansal et al., 2021; Prabhudesai et al., 2023); but this relies on well-calibrated confidence scores, which are often difficult to achieve—models can be wrong with high confidence.
@@ 2.2. Alternative Forms of AI Decision Support
Beyond abstract concepts, concrete examples for supporting decisions without relying on end-to-end recommendations are often found in healthcare. Lindvall et al. (Lindvall et al., 2021) designed a system for tumor assessment that navigates pathologists to potentially tumorous image regions for them to review, without revealing whether or how confidently the model classifies the pixels as tumor. Crucially, the classification threshold was not set to optimize accuracy, but sensitivity. Consequently, if pathologists did not find a tumor in any of the suggested regions, they could be relatively sure that the rest of the image also contains no cancer. Zhang et al. (Zhang et al., 2023b) studied the case of sepsis diagnosis, where an existing AI tool only addresses the final stage of the decision-making by offering a sepsis risk score and sending alerts above a certain threshold. The authors proposed a redesign where the system suggests lab tests that would help to reduce uncertainty about patients’ future conditions.
In the context of aviation, Zhang et al. explored how to support diversions decisions (Zhang et al., 2023a). Their system does not recommend airports, but continuously provides unobtrusive local hints about potential limitations at the surrounding airports. The system provides this support of pilots’ SA also during normal flight when there is no sign of an emergency yet.
@@ 3.1. System Design and Variants
Table with four versions of the diversion assistance.
@@
Screenshots of the system.
@@ 3.2. Apparatus
Schematic overview of study apparatus.
@@ 4. Methods
Schematic overview of the study parts.
@@ 4.1.1. Scenario 1: Single engine failure
Screenshot of single engine failure scenario.
@@ 4.1.2. Scenario 2: Passenger medical emergency
Screenshot of passenger medical emergency scenario.
@@ 4.1.3. Scenario 3: Airport closure
Screenshot of airport closure scenario.
This last scenario was based on the personal experience of one participant in Zhang et al.’s study (Zhang et al., 2023a). The flight was from Sharm El Sheikh, Egypt, to Hamburg, Germany (Figure 8). About an hour before landing, participants were given the information that the destination had closed down, again via a popup message: “Due to a power outage, EDDH [^13] The ICAO (International Civil Aviation Organization) airport code for Hamburg. had to close completely. It will remain closed for at least the next few hours.”
@@ 5.1.2. Differences between scenarios and support paradigms
A table of the different usage patterns.
@@ 5.2. RQ2: Overreliance
Point plot of overreliance occurrences.
@@ 6.1. There Is More to Decision Support than Giving Recommendations and Explanations
The common, but often tacit design goal for AI decision support is to solve the task for users, creating a redundant rather than complementary role to the human. While it is popular to speak of “human-AI collaboration” in AI-assisted decision-making research (Schmidt and Biessmann, 2020; Hemmer et al., 2023), users may perceive this recommendation-centric support more as “human-AI competition” (Zhang et al., 2023b). Some of our participants expressed similar concerns regarding recommendations.
Like previous studies involving real-world tasks with experts, our participants’ views suggest a shift in the design goal from solving the entire decision task for users to addressing their primary pain points. For the tumor assessment use case studied by Lindvall et al. (Lindvall et al., 2021), pathologists’ biggest challenge was to find small tumorous regions in huge images. In the case of sepsis diagnosis studied by Zhang et al. (Zhang et al., 2023b), physicians wanted to know which lab tests to order, when. In our case of diversions, pilots would benefit the most from a quick overview of a large amount of relevant information. This shift in design goal enables much more diverse uses for AI than merely giving end-to-end recommendations, as exemplified by the various suggestions by our participants.
The challenge shifts from giving best possible recommendations and explaining them, to serving pilots’ momentary information needs. Transparency is consequently not only—or maybe not even primarily—required to calibrate users’ reliance on the recommendation, as is the current focus in AI-assisted decision-making research (Fok and Weld, 2023). Rather, the role of transparency is to make visible how well the information fits pilots’ current intention.
@@ 6.3. From Recommendation-Centric to Process-Oriented Support
Bringing together the discussed aspects, we propose process-oriented decision support as a promising alternative framework to design AI decision support that is not centered around end-to-end recommendations. Figure 12 shows how process-oriented support compares to recommendation-centric support, and how it applies to the diversion use case through both continuous support and participants’ design suggestions. We consider the framework to be applicable not only for diversion assistance but more generally for AI support of high-stakes decisions. We argue that process-oriented support is particularly useful for complex decisions where human expertise and context knowledge is crucial, but hard to combine with end-to-end recommendations, such as in healthcare (Zhang et al., 2023b), social work (Kawakami et al., 2022b), or sales (Blomberg et al., 2018). The key difference to recommendation-centric support is the shift from trying to solve the task for users through end-to-end recommendations, to helping users to solve the task by addressing the challenges in their decision-making process. These challenges have to be determined for each application through user research. In our diversion use case, pilots’ main challenge was to gather the information they need to make a decision.
@@
Lorem.
@@
Second, the role of transparency changes. In recommendation-centric support, transparency aids appropriate reliance on recommendations. In process-oriented support, transparency facilitates appropriation (Dix, 2007) by helping users understand how well the system functionality aligns with their momentary intentions and needs. The importance of appropriation in AI-assisted decision-making has been discussed by several authors recently (Zhang et al., 2023a; Cabitza et al., 2021; Ehsan et al., 2021). It differs from reliance calibration in at least two aspects. One is that transparency for appropriation is targeting the more granular level of intermediary decision-making steps, rather than the end of the process. This leads to more nuanced behavior than a simple accept-reject dichotomy, like partial reliance and negotiation patterns, as observed by Sivaraman et al. (Sivaraman et al., 2023) and with our participants as well. The other difference is that appropriate reliance concerns the correctness of the AI, while appropriation is about the user’s intention. As was evident from our participants’ requests, this does not necessarily involve explaining how the algorithm works.

### v3

@@ Abstract.
copyright: acmlicensedjournal: PACMHCIjournalyear: 2024journalvolume: 8journalnumber: CSCW2article: 485publicationmonth: 11doi: 10.1145/3687024ccs: Information systems Decision support systemsccs: Human-centered computing Interaction paradigms
@@
Schematic diagram of both support concepts. There two main sections that fade into each other: normal flight to the left, and abnormal situation to the right. In the middle, there are six blocks connected by arrows, which are from left to right: hypothetical options and hypothetical scenarios for normal flight, and facts, options, risks and benefits, and decision for abnormal situation. Above these blocks, there is a thick arrow labeled recommendation-centric support (backward). It points from right to left throughout the abnormal situation section. Smaller arrows split out of it down to the options, risks and benefits, and decision blocks. The one pointing into the decision block is labeled 1) recommend decision. The two other arrows are labeled 2) help to evaluate the AI recommendation. Below the central block is another thick arrow, pointing from left to right throughout the entire figure. One smaller arrow splits out and points into the hypothetical options block, labeled 1) help to evaluate the surrounding options. Two further smaller arrow point into the options and risks and benefits blocks, labeled 2) help to evaluate the surrounding options. A last smaller arrow, which has a slightly faded color, points into the decision block, labeled 3) recommend decision based on evaluation.
@@ 1. Introduction
Yet, in spite of good machine performance, real-world adoption of AI is often difficult (Schlögl et al., 2019; Varghese, 2020). In controlled studies, failure to achieve complementary performance is often observed (Bansal et al., 2021; Liu et al., 2021; Fok and Weld, 2024), i.e., the combination of human and AI performs worse than one of them alone. This is the result of inappropriate reliance, including both overreliance (human relies on AI even when it is disadvantageous to do so) (Bansal et al., 2021; Liu et al., 2021; Jacobs et al., 2021b; Schmidt and Biessmann, 2020; Bussone et al., 2015; Buçinca et al., 2021; Lai and Tan, 2019; Poursabzi-Sangdeh et al., 2021; Wang and Yin, 2021; Green and Chen, 2019) and underreliance (human rejects AI even when it would be beneficial to rely on it) (Dietvorst et al., 2015; Cheng and Chouldechova, 2023; Castelo et al., 2019; Prahl and Van Swol, 2017).
@@
In practice, AI support often turns out less useful to decision makers than imagined (Yang et al., 2016; Blomberg et al., 2018; Kawakami et al., 2022a). A growing number of formative studies on real-world tasks with domain experts (Kawakami et al., 2022a, b; Jacobs et al., 2021a; Sivaraman et al., 2023; Bach et al., 2023; Burgess et al., 2023; Kaltenhauser et al., 2020; Yang et al., 2016; Zhang et al., 2024) try to understand what hinders effective use of AI decision support. A frequent problem is that AI decision support is usually designed to be recommendation-centric, where the primary functionality of the system is to give end-to-end decision recommendations, i.e., the system suggests a possible end result straight from its input data. By directly jumping to the end result, these systems only support the very end of the decision-making (Buçinca et al., 2022), ignoring the entire process leading up to the decision (Yang et al., 2016; Zhang et al., 2024, 2023).
@@
While the limitations of recommendation-centric support are becoming increasingly more apparent, there is a notable lack of effective solutions to the identified problems. Current research is mostly limited to proposing alternatives for recommendation-centric support on a conceptual level (Zhang et al., 2021; Buçinca et al., 2022; Cabitza et al., 2021; Miller, 2023; Koon, 2022), but concrete examples are few and far between, with most examples stemming from healthcare (Cai et al., 2019; van Berkel et al., 2021; Yang et al., 2019; Lindvall et al., 2021; Gu et al., 2023; Zhang et al., 2024). Even fewer works evaluate these alternative support paradigms in comparison to recommendation-centric decision support.
@@
One of the few studies exploring alternative decision support roles for AI is Zhang et al.’s work on diversion assistance (Zhang et al., 2023). Their key insight was that diversion decisions are not a point, but a process in which pilots take proactive actions. Even during normal flight, pilots constantly establish situation awareness (SA) and prepare a valid plan B, should an emergency occur. Continuously supporting these proactive actions via unobtrusive hints was found to be a promising support role for AI, but it remains unclear how effective it would be in practice, especially compared to the established recommendation-centric paradigm.
@@ 2.1. Recommendation-Centric Decision Support
The most common strategy to help decision makers work better with recommendation-centric AI is to add explanations of how the model works (Wang and Yin, 2021; Bussone et al., 2015; Bansal et al., 2021; Zhang et al., 2020; Lai and Tan, 2019) and other model information, such as model confidence (Zhang et al., 2020; McGuirl and Sarter, 2006; Rechkemmer and Yin, 2022) or the stated model accuracy (He et al., 2023; Rechkemmer and Yin, 2022). The goal is to help people to rely appropriately on AI recommendations (Schemmer et al., 2023) by giving cues about when they may be beneficial or detrimental to rely on. Results have been mixed so far, as especially explanations are prone to induce blind trust (Bansal et al., 2021; Eiband et al., 2019) and hence overreliance, even among domain experts (Bussone et al., 2015; Jacobs et al., 2021b). Recently, Vasconcelos et al. (Vasconcelos et al., 2023) have shown that explanations can reduce overreliance under certain conditions, but it is questionable how often these conditions are valid in real applications (Fok and Weld, 2024). Communicating model confidence appears more promising, as it has repeatedly been shown to improve appropriate reliance (Zhang et al., 2020; McGuirl and Sarter, 2006; Bansal et al., 2021; Prabhudesai et al., 2023); but this relies on well-calibrated confidence scores, which are often difficult to achieve—models can be wrong with high confidence.
@@ 2.2. Alternative Forms of AI Decision Support
Beyond abstract concepts, concrete examples for supporting decisions without relying on end-to-end recommendations are often found in healthcare. Lindvall et al. (Lindvall et al., 2021) designed a system for tumor assessment that navigates pathologists to potentially tumorous image regions for them to review, without revealing whether or how confidently the model classifies the pixels as tumor. Crucially, the classification threshold was not set to optimize accuracy, but sensitivity. Consequently, if pathologists did not find a tumor in any of the suggested regions, they could be relatively sure that the rest of the image also contains no cancer. Zhang et al. (Zhang et al., 2024) studied the case of sepsis diagnosis, where an existing AI tool only addresses the final stage of the decision-making by offering a sepsis risk score and sending alerts above a certain threshold. The authors proposed a redesign where the system suggests lab tests that would help to reduce uncertainty about patients’ future conditions.
In the context of aviation, Zhang et al. explored how to support diversions decisions (Zhang et al., 2023). Their system does not recommend airports, but continuously provides unobtrusive local hints about potential limitations at the surrounding airports. The system provides this support of pilots’ SA also during normal flight when there is no sign of an emergency yet.
@@ 3.1. System Design and Variants
Table with four versions of the diversion assistance. The rows are the four versions, the columns are normal flight, emergency occurs, and select diversion option. The table shows smaller versions of the screeshots shown in Figure 3. The normal flight column is populated with screenshots of the local hints for Cont and Rec+Cont, and empty for the other two rows. The emergency occurs column holds screenshots for select emergency type for all rows except for Baseline. The select diversion option column holds for the Rec and Rec+Cont rows screenshots for recommendations and in brackets for edit recommendation criteria. For the Cont row, there is screenshot for adjusted local hints. For the Baseline row, there is a screenshot for table without AI elements.
@@
Six numbered screenshots of the system, matching the numbers in the caption, arranged in a grid with two columns and three rows. Screenshots 1 (recommendations), 3 (local hints during normal flight), and 4 (adjusted local hints during emergency), magnify in circles the respective parts of the screenshots.
@@ 3.2. Apparatus
Three main blocks. Top left: diversion assistance front end, with a tablet and logos of Framer and React. Top right: flight simulator, with a monitor showing the A320 cockpit in X-Plane, the X-Plane 11 logo, and depictions of the FCU, the MCDU, throttle, and sidestick. At the bottom: block with database and Python back end, with arrow pointing from the former to the latter. The arrow is labeled hard-coded scenario data, with bullets reading weather, runways, airport infrastructure, NOTAMs. An arrow points from the bottom block to the front end block, labeled Websocket connection. Another arrow points from the bottom block to the flight simulator block, labeled timed emergency triggers. Another arrow points the opposite direction, labeled live fight data, with bullets reading aircraft position, speed, heading and fuel flow; airport positions.
@@ 4. Methods
Pipeline starting with initial scenario outlines on the left, followed by a right-pointing arrow, followed by focus group (N=4), followed by another right-pointing arrow, followed by pilot test (N=2). Below pilot test, there is a block for system design, with an arrow pointing up to pilot test. To the right of pilot test is another right-pointing arrow, followed by final scenario and system design, followed by yet another right-pointing arrow and main study (N=32). Below main study are three bullet points, reading 4 conditions (system variants), between-subject; 3 scenarios with think-alouds; exit interview.
@@ 4.1.1. Scenario 1: Single engine failure
Screenshot of single engine failure scenario, showing on a map with a line the flight route, and with colored crosses departure (Geneva) and destination (London) as well as notable en-route airports (Paris, Lyon), and the place where the event (Engine 1 fails) happened.
@@ 4.1.2. Scenario 2: Passenger medical emergency
Screenshot of passenger medical emergency scenario, showing on a map with a line the flight route, and with colored crosses departure (Tenerife) and destination (Munich) as well as notable en-route airports (Casablanca, Faro), and the place where the event (passenger has heart attack) happened.
@@ 4.1.3. Scenario 3: Airport closure
Screenshot of airport closure scenario, showing on a map with a line the flight route, and with colored crosses departure (Sharm El Sheikh) and destination (Hamburg) as well as notable en-route airports (Berlin, Hanover, Düsseldorf), and the place where the event (information that Hamburg is closed) happened.
This last scenario was based on the personal experience of one participant in Zhang et al.’s study (Zhang et al., 2023). The flight was from Sharm El Sheikh, Egypt, to Hamburg, Germany (Figure 8). About an hour before landing, participants were given the information that the destination had closed down, again via a popup message: “Due to a power outage, EDDH [^13] The ICAO (International Civil Aviation Organization) airport code for Hamburg. had to close completely. It will remain closed for at least the next few hours.”
@@ 5.1.2. Differences between scenarios and support paradigms
A very complex table of the different usage patterns, with shades of blue to visualize the frequency of pattern occurances. Columns hold the patterns, categorized by options considered and decision strategy. The latter is further split into general strategies and AI usage strategies. The latter is again further split into normal flight mode, highlights, and recommendations. The rows hold the four conditions, each three time for the three scenarios.
@@ 5.2. RQ2: Overreliance
Point plot of overreliance occurrences, with conditions on y-axis and overreliance probability on x-axis. The probabilities are: Baseline 0.125, Rec+Cont 0.375, Cont 0.5, Rec 0.75. The error bars are very long, spanning around 0.4 for each condition.
@@ 5.3. RQ3: Decision Time
[…]
Line plot with error bars. Decision time in seconds on y-axis, from 0 to 400, and the four conditions on x-axis. Baseline has clearly the highest decision time (around 275 seconds), followed by Rec, Cont, and Rec+Cont, in descending order, which are all relatively close to each other (all between 150 and 190 seconds).
@@
[…]
Line plot with error bars. Decision time in seconds on y-axis, from 0 to 200, and the four conditions on x-axis. Rec+Cont has clearly the lowest decision time (around 60 seconds), followed by Cont, Baseline, and Rec in ascending order, which are all relatively close to each other (all between 125 and 150 seconds).
@@
[…]
Line plot with error bars. Decision time in seconds on y-axis, from 0 to 300, and the four conditions on x-axis. Baseline has by a small margin the highest decision time (around 270 seconds). Rec, Cont, Rec+Cont, in ascending order, are close to each other (all around 225 seconds).
@@ 6.1. There Is More to Decision Support than Giving Recommendations and Explanations
The common, but often tacit design goal for AI decision support is to solve the task for users, creating a redundant rather than complementary role to the human. While it is popular to speak of “human-AI collaboration” in AI-assisted decision-making research (Schmidt and Biessmann, 2020; Hemmer et al., 2023), users may perceive this recommendation-centric support more as “human-AI competition” (Zhang et al., 2024). Some of our participants expressed similar concerns regarding recommendations.
Like previous studies involving real-world tasks with experts, our participants’ views suggest a shift in the design goal from solving the entire decision task for users to addressing their primary pain points. For the tumor assessment use case studied by Lindvall et al. (Lindvall et al., 2021), pathologists’ biggest challenge was to find small tumorous regions in huge images. In the case of sepsis diagnosis studied by Zhang et al. (Zhang et al., 2024), physicians wanted to know which lab tests to order, when. In our case of diversions, pilots would benefit the most from a quick overview of a large amount of relevant information. This shift in design goal enables much more diverse uses for AI than merely giving end-to-end recommendations, as exemplified by the various suggestions by our participants.
The challenge shifts from giving best possible recommendations and explaining them, to serving pilots’ momentary information needs. Transparency is consequently not only—or maybe not even primarily—required to calibrate users’ reliance on the recommendation, as is the current focus in AI-assisted decision-making research (Fok and Weld, 2024). Rather, the role of transparency is to make visible how well the information fits pilots’ current intention.
@@ 6.3. From Recommendation-Centric to Process-Oriented Support
Bringing together the discussed aspects, we propose process-oriented decision support as a promising alternative framework to design AI decision support that is not centered around end-to-end recommendations. Figure 12 shows how process-oriented support compares to recommendation-centric support, and how it applies to the diversion use case through both continuous support and participants’ design suggestions. We consider the framework to be applicable not only for diversion assistance but more generally for AI support of high-stakes decisions. We argue that process-oriented support is particularly useful for complex decisions where human expertise and context knowledge is crucial, but hard to combine with end-to-end recommendations, such as in healthcare (Zhang et al., 2024), social work (Kawakami et al., 2022b), or sales (Blomberg et al., 2018). The key difference to recommendation-centric support is the shift from trying to solve the task for users through end-to-end recommendations, to helping users to solve the task by addressing the challenges in their decision-making process. These challenges have to be determined for each application through user research. In our diversion use case, pilots’ main challenge was to gather the information they need to make a decision.
@@
Schematic overview of recommendation-centric decision support on top, and process-oriented decision support on the bottom. In the former, the user is outside of the decision-making process, while in the latter, the user is within the decision-making process.
@@
Second, the role of transparency changes. In recommendation-centric support, transparency aids appropriate reliance on recommendations. In process-oriented support, transparency facilitates appropriation (Dix, 2007) by helping users understand how well the system functionality aligns with their momentary intentions and needs. The importance of appropriation in AI-assisted decision-making has been discussed by several authors recently (Zhang et al., 2023; Cabitza et al., 2021; Ehsan et al., 2021). It differs from reliance calibration in at least two aspects. One is that transparency for appropriation is targeting the more granular level of intermediary decision-making steps, rather than the end of the process. This leads to more nuanced behavior than a simple accept-reject dichotomy, like partial reliance and negotiation patterns, as observed by Sivaraman et al. (Sivaraman et al., 2023) and with our participants as well. The other difference is that appropriate reliance concerns the correctness of the AI, while appropriation is about the user’s intention. As was evident from our participants’ requests, this does not necessarily involve explaining how the algorithm works.